How do I solve the following equations
5/6x=10
2/3x= -3/5

Answer by rfer(16322)

5/6x-10
x=10/(5/6)
x=10*6/5
x=12
-------------
2/3x=-3/5
x=(-3/5)/(2/3)
x=-3/5*3/2
x=-9/10